# Conditional regression for the Nonlinear Single-Variable Model

### v1

@@ Abstract
Several statistical models for regression of a function $F$ on $\mathbb{R}^{d}$ without the statistical and computational curse of dimensionality exist, for example by imposing and exploiting geometric assumptions on the distribution of the data (e.g. that its support is low-dimensional), or strong smoothness assumptions on $F$, or a special structure $F$. Among the latter, compositional models assume $F=f\circ g$ with $g$ mapping to $\mathbb{R}^{r}$ with $r\ll d$, have been studied, and include classical single- and multi-index models and recent works on neural networks. While the case where $g$ is linear is rather well-understood, much less is known when $g$ is nonlinear, and in particular for which $g$’s the curse of dimensionality in estimating $F$, or both $f$ and $g$, may be circumvented. In this paper, we consider a model $F(X):=f(\Pi_{\gamma}X)$ where $\Pi_{\gamma}:\mathbb{R}^{d}\to[0,\textrm{len}_{\gamma}]$ is the closest-point projection onto the parameter of a regular curve $\gamma:[0,\textrm{len}_{\gamma}]\to\mathbb{R}^{d}$ and $f:[0,\textrm{len}_{\gamma}]\to\mathbb{R}^{1}$. The input data $X$ is not low-dimensional, far from $\gamma$, conditioned on $\Pi_{\gamma}(X)$ being well-defined. The distribution of the data, $\gamma$ and $f$ are unknown. This model is a natural nonlinear generalization of the single-index model, which corresponds to $\gamma$ being a line. We propose a nonparametric estimator, based on conditional regression, and show that under suitable assumptions, the strongest of which being that $f$ is coarsely monotone, it can achieve the one-dimensional optimal min-max rate for non-parametric regression, up to the level of noise in the observations, and be constructed in time $\mathcal{O}(d^{2}n\log n)$. All the constants in the learning bounds, in the minimal number of samples required for our bounds to hold, and in the computational complexity are at most low-order polynomials in $d$.
Keywords: Nonparametric regression; compositional models; single-index model.
@@ 1 Introduction
In the general nonparametric and distribution-free setup where we only know that $F\in\mathcal{C}^{s}(\mathbb{R}^{d})$ is Hölder continuous with exponent $s>0$ and, say, compacted supported, the min-max nonparametric rate for estimating $F$ in $L^{2}(\rho)$ is $n^{-\frac{s}{2s+d}}$, see (Binev et al., 2023; Györfi et al., 2002) and references therein. Kernel estimators attain this learning rate with properly chosen bandwidth and kernel, as so do a variety of other well-understood estimators based on Fourier or multiscale decompositions, see (Györfi et al., 2002; Binev et al., 2005, 2007) and the numerous references therein. This rate deteriorates dramatically as the dimension $d$ of the ambient space increases: this is an instance of the curse of dimensionality. As no estimator can achieve a faster learning rate, in the min-max sense, for all such functions, and yet in many applications $d$ is very large, it is of interest to consider special classes of regression problems where the curse of dimensionality may be avoided.
@@
Intrinsically low-dimensional models. In one direction, one may make geometric assumptions about the distribution $\rho_{X}$ of the input data $X_{i}$, for example that $\rho_{X}$ is supported on a low-dimensional manifold $\mathcal{M}$, say of dimension $r\ll d$, while $F$ is, say, a generic $s$-Hölder function on $\mathcal{M}$. In these settings, estimators exist that converge at an optimal rate concerning the intrinsic dimension $r$, see (Bickel and Li, 2007; Kpotufe, 2011; Kpotufe and Garg, 2013; Liao et al., 2016, 2022; Liu et al., 2024a), with at least some of them with associated fast algorithms, for example the estimators in (Liao et al., 2016, 2022) can be constructed in time $\mathcal{O}(C^{r}d^{s}n\log n)$ and they achieve, adaptively, min-max optimal rate on large function spaces with unknown regularity possibly varying across locations and scales.
@@
Functions with high degree of regularity. In a different direction, one may consider function classes with smaller complexity (e.g., as measured by metric entropy) than $s$-Hölder functions on $\mathbb{R}^{d}$. A straightforward but highly limiting assumption is that the Hölder exponent $s$ is proportional to $d$ so that the min-max rate above is independent of $d$. Larger function classes are obtained by imposing integrability of the Fourier transform (Barron, 1993), or strong mixed-smoothness (Strömberg, 1998), for example by using highly anisotropic Besov spaces. These choices however still require a degree of smoothness increasing with $d$ in order to avoid the curse of dimensionality. For example in these anisotropic Besov spaces, functions need to have $\mathcal{O}(1)$ axis-oriented directions in which they have “regular, $\mathcal{O}(1)$ smoothness”, and have “$\mathcal{O}(d)$ smoothness” in all other directions. In the context of deep neural networks, Suzuki and Nitanda (2024) demonstrate that functions in this class can be well-approximated well, with suitable architectures, in a way that avoids the curse of dimensionality if the degree of anisotropic smoothness is sufficiently large, while leaving open the aspects of learning and optimization.
@@
having some Hölder regularity, and $r\leq d$. For general function compositions, (Juditsky et al., 2009) proves that in some situations, when $g$ is sufficiently smooth, there is an improved min-max nonparametric rate of estimation of $F=f\circ g$, but still subject to the curse of dimensionality. Recent works combine anisotropic smoothness with composability, especially in the context of approximators or estimators constructed with deep neural networks. In this direction, (Schmidt-Hieber, 2020; Shen et al., 2021) consider spaces of functions that are compositions of low-dimensional functions with anisotropic smoothness conditions, studying the dependence of approximability and learning risk on the dimensionality of the spaces involved and the smoothness of the corresponding functions. Unfortunately the inner-most such function is defined on the original high-dimensional space $\mathbb{R}^{d}$, so in general these results do not circumvent the curse of dimensionality, unless again the (anisotropic, in general) regularity of such function scales with $d$. Another direction is to design different types of neural networks that exploit nonlinear compositions, see (Lai and Shen, 2021; Liu et al., 2024b). Most of these results only address whether a function can well-approximated by a neural network, and do not cover the learning problem nor address computational aspects, in particular whether some optimization algorithm can find (efficiently?) the parameters of the desired estimator. Another aspect that is often left unaddressed is the dependency of constants on the ambient dimension, which is unfortunately often exponential: see Shamir (2020) for a discussion of some of these aspects, and Shen et al. (2021) for bounds on such constants in some of these cases. In the parametric setting, models where $g$ is a polynomial and $f$ is a function in a known finite-dimensional space have been considered; for example Wang et al. (2024); Lee et al. (2024) (and related work referenced therein) consider special classes of polynomials $g$ for which a customized layer-by-layer training of a neural network leads to estimators of $F$ that are not cursed by the dimensionality.
@@
Single- and multi-index models. Classical examples of structural assumptions based on function composition for which the curse of dimensionality can provably be avoided include single- and multi-index models, as well as generalized linear models (Stone, 1982; Hastie and Tibshirani, 2014; Breiman and Friedman, 1985; Horowitz and Mammen, 2007). In the case when $g=G:\mathbb{R}^{d}\to\mathbb{R}^{r}$ is a linear operator, this is called the multi-index model, and it implies that the function $F\in\mathcal{C}^{s}(\mathbb{R}^{d})$ only depends on a small number $r$ of linear features: $F(x)=f(Gx)$ for some link function $f\in\mathcal{C}^{s}(\mathbb{R}^{r})$, matrix $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{r}$, and the projection of $X$ on the range of $G$ is sufficient for regression. The particular case $r=1$ is called the single-index model, where the function $F$ has the structure $F(x)=f(\langle v,x\rangle)$ for some unknown index vector $v\in\mathbb{S}^{d-1}$ and unknown link function $f\in\mathcal{C}^{s}(\mathbb{R}^{1})$. These models have been intensively studied: Stone (1982) conjectured that the min-max rate for regression of single-index models is $n^{-\frac{s}{2s+1}}$ (resp. $n^{-\frac{s}{2s+r}}$ for multi-index models), coinciding with the one-dimensional min-max nonparametric rate, thereby escaping the curse of dimensionality. This rate is achieved with kernel estimators in (Hardle and Stoker, 1989, Theorem 3.3) and (Horowitz, 1998, Section 2.5), and the index $v$ can be estimated at the parametric rate $\mathcal{O}(n^{-\frac{1}{2}})$ under suitable assumptions. The existence of an estimator converging at rate $n^{-\frac{s}{2s+1}}$ is shown in (Györfi et al., 2002, Corollary 22.1), and (Gaïffas and Lecué, 2007, Theorem 2) demonstrates that $n^{-\frac{s}{2s+1}}$ is indeed the min-max rate. These models have also received attention recently both for their connections to “feature learning” in neural networks, see for example Lee et al. (2024) and references therein for learning single-index models with neural networks (trained in a suitable layer-wise fashion), as well as Radhakrishnan et al. (2024), and references therein, showing that deep neural networks appear to exploit low-dimensional linear subspaces for classification and prediction.
@@
Several methods for estimating the index vector $v$ (or the matrix $G$) and $f$ were developed over the years, with varying tradeoffs between the assumptions needed to produce optimal estimators and the computational cost for constructing them. One category includes semiparametric methods based on maximum likelihood estimation (Ichimura, 1993; Hardle et al., 1993; Delecroix et al., 1997; Delecroix and Hristache, 1999; Delecroix et al., 2006; R. J. Carroll and Wand, 1997) and M-estimators that produce $\sqrt{n}$-consistent index estimates under general assumptions, but with computationally demanding implementations, relying on sensitive bandwidth selections for kernel smoothing and on high-dimensional joint optimization of $f$ and $v$. Another category includes direct methods: for example Average Derivative Estimation (Stoker, 1986; Hardle and Stoker, 1989) estimates the index vector $v$ by exploiting its proportionality to $\nabla F$. Early implementations of this idea suffer from the curse of dimensionality due to the use of kernel estimation for the gradient. The technique of (Hristache et al., 2001) uses an iterative scheme to gradually adapt an elongated neighbourhood window. Gradient-based methods (stochastic and non-stochastic) have been studied in the context of neural networks (Lee et al., 2024).
@@
A category of techniques particularly relevant to the present work includes conditional, or inverse regression methods, which derive their estimators from statistics of the conditional distribution of the explanatory variable $X$ given observations of the (noisy) response variable $Y$. Prominent examples include sliced inverse regression (Duan and Li, 1991; Li, 1991), sliced average variance estimation (Cook, 2000), simple contour regression (Li et al., 2005; Coudret et al., 2013), with its analysis (in the multi-index case) in the work (Li and Wang, 2007), which yields an estimator with near-optimal rates for the multi-index model. Conditional methods partition the range of the response variable $Y$ into small intervals and consider their pre-images that are, when $f$ is monotone, slices with the thinnest side along the index direction $v$. They are often straightforward in implementation, consisting of the computation of conditional empirical moments, or other statistics related to the level sets of $f$, and can have only one parameter to tune: the width of slices. Unfortunately, several of the works above, including (Hristache et al., 2001; Li and Wang, 2007), while obtaining asymptotic learning rates that avoid the curse of dimensionality and are in some cases min-max optimal, require a minimal number of samples that is exponential in the ambient dimension $d$ (either explicitly, or through constants exponential in $d$ in the bounds).
@@
Besides the statistical view that the single- and multi-index model do not entail a statistical cost cursed by the ambient dimension, another critical problem is to find algorithms with reasonable computational cost to implement a statistically optimal estimator for these models. For example, even in the case of single-index models, $M$-estimators typically required high-dimensional non-convex optimization in $v$ and $f$; methods based on optimizing over $v$, even when combined with random sampling, may scale exponentially in $d$ due to the need of obtaining points inside a narrow cone around the unknown $v$. Conditional methods can often be implemented in a computationally efficient way, running in time $\mathcal{O}(C_{d}n)$ (up to $\log n$ factors, and with the exception of contour regression methods, which scale quadratically in $n$), with a constant $C_{d}$ a low-order polynomial in $d$. Lanteri et al. (2022) discuss in detail these techniques for the single-index model, and introduce of a variation, called Smallest Vector Regression, which enjoys optimal statistical guarantees up to log factors, provides a theoretically optimal choice for selecting of the slice width parameter (a crucial parameter whose choice is often not discussed (Coudret et al., 2013, p. 75)), and is amenable to a computationally-efficient implementation with cost $\mathcal{O}(d^{2}n\log n)$, all without the curse of dimensionality in the exponents and in the constants, for the rate, the minimum sample requirements and the computational cost.
@@
In conclusion, while the situation is rather well-understood for compositional models $F=f\circ g$ with $g$ linear, much is open when $g$ is nonlinear, in which case it is not clear when the compositional structure allows one to circumvent the curse of dimensionality, except in cases where $g$ has high regularity (global or anisotropic), increasing with $d$, or belong to certain families of polynomials. We propose a simple model that, perhaps surprisingly, without such strong regularity assumptions, still is amenable to nonparametric estimators that avoid the curse of dimensionality, both statistically and computationally.

### v2

@@ Abstract
Regressing a function $F$ on $\mathbb{R}^{d}$ without the statistical and computational curse of dimensionality requires special statistical models, for example that impose geometric assumptions on the distribution of the data (e.g., that its support is low-dimensional), or strong smoothness assumptions on $F$, or a special structure $F$. Among the latter, compositional models $F=f\circ g$ with $g$ mapping to $\mathbb{R}^{r}$ with $r\ll d$ include classical single- and multi-index models, as well as neural networks. While the case where $g$ is linear is well-understood, less is known when $g$ is nonlinear, and in particular for which $g$’s the curse of dimensionality in estimating $F$, or both $f$ and $g$, may be circumvented. Here we consider a model $F(X):=f(\Pi_{\gamma}X)$ where $\Pi_{\gamma}:\mathbb{R}^{d}\to[0,\textrm{len}_{\gamma}]$ is the closest-point projection onto the parameter of a regular curve $\gamma:[0,\textrm{len}_{\gamma}]\to\mathbb{R}^{d}$, and $f:[0,\textrm{len}_{\gamma}]\to\mathbb{R}^{1}$. The input data $X$ is not low-dimensional: it can be as far from $\gamma$ as the condition that $\Pi_{\gamma}(X)$ is well-defined allows. The distribution $X$, the curve $\gamma$ and the function $f$ are all unknown. This model is a natural nonlinear generalization of the single-index model, corresponding to $\gamma$ being a line. We propose a nonparametric estimator, based on conditional regression, that under suitable assumptions, the strongest of which being that $f$ is coarsely monotone, achieves, up to log factors, the one-dimensional optimal min-max rate for non-parametric regression, up to the level of noise in the observations, and be constructed in time $\mathcal{O}(d^{2}n\log n)$. All the constants in the learning bounds, in the minimal number of samples required for our bounds to hold, and in the computational complexity are at most low-order polynomials in $d$.
Keywords: High-dimensional regression; nonparametric regression; compositional models; single-index model.
@@ 1 Introduction
In the general nonparametric and distribution-free setup where we only know that $F\in\mathcal{C}^{s}(\mathbb{R}^{d})$ is Hölder continuous with exponent $s>0$ and, say, compactly supported, the min-max nonparametric rate for estimating $F$ in $L^{2}(\rho_{X})$ is $n^{-\frac{s}{2s+d}}$, see (Binev et al., 2023; Györfi et al., 2002) and references therein. Kernel estimators attain this learning rate with properly chosen bandwidth and kernel, as so do a variety of other well-understood estimators based on Fourier or multiscale decompositions, see (Györfi et al., 2002; Binev et al., 2005, 2007) and the numerous references therein. This rate deteriorates dramatically as the dimension $d$ of the ambient space increases: this is an instance of the curse of dimensionality. As no estimator can achieve a faster learning rate, in the min-max sense, for all such functions, and yet in many applications $d$ is very large, it is of interest to consider special classes of regression problems where the curse of dimensionality may be avoided.
In this work, we introduce and construct efficient estimators for the model where $F(x)=f(\Pi_{\gamma}x)$, where $\gamma$ is a curve in $\mathbb{R}^{d}$, and $\Pi_{\gamma}$ maps $x\in\mathbb{R}^{d}$ to the closest point to $x$ that belongs to $\gamma$. The input data $x$ is sampled from a distribution supported around $\gamma$, and it is therefore not low-dimensional. With the curve $\gamma$, the distribution of the input data, and the function $f$ all unknown, we construct an estimator for $F$, using conditional inverse regression techniques, that converges at the near min-max one-dimensional rate for regression (up to log factors), and that can constructed in $\mathcal{O}(d^{2}n\log n)$, where $n$ is the number of samples. It therefore avoids both the statistical and the computational curse of dimensionality. The model is detailed in the next section, together with an informal statement of the main theorem. Before we delve into details, we provide some context and motivation for this work, and in particular discuss statistical models for regression that have been introduced and studied in order to avoid the curse of dimensionality.
@@
Intrinsically low-dimensional models. In one direction, one may make geometric assumptions about the distribution $\rho_{X}$ of the input data $X_{i}$, for example that $\rho_{X}$ is supported on a low-dimensional manifold $\mathcal{M}$, say of dimension $r\ll d$, while $F$ is, say, a generic $s$-Hölder function on $\mathcal{M}$. In these settings, estimators exist that converge at an optimal rate with respect to the intrinsic dimension $r$, see (Bickel and Li, 2007; Kpotufe, 2011; Kpotufe and Garg, 2013; Liao et al., 2016, 2022; Liu et al., 2024a), with at least some of them with associated fast algorithms, for example the estimators in (Liao et al., 2016, 2022) can be constructed in time $\mathcal{O}(C^{r}d^{s}n\log n)$ and they achieve, adaptively, min-max optimal (up to log factors) rate on a large family of function spaces with unknown regularity, possibly varying across locations and scales.
@@
Functions with high degree of regularity. In a different direction, one may consider function classes with smaller complexity (e.g., as measured by metric entropy) than $s$-Hölder functions on $\mathbb{R}^{d}$. One straightforward but highly limiting assumption is that the Hölder exponent $s$ is proportional to $d$ so that the min-max rate above is independent of $d$. Larger function classes are obtained by either imposing strong mixed-smoothness (Strömberg, 1998) such as highly anisotropic Besov spaces (where functions need to have $\mathcal{O}(1)$ axis-oriented directions in which they have “regular, $\mathcal{O}(1)$ smoothness”, and have “$\mathcal{O}(d)$ smoothness” in all other directions), or imposing integrability of the Fourier transform (Barron, 1993). In the context of deep neural networks, Suzuki and Nitanda (2024) demonstrate that functions in this class can be approximated well, with suitable architectures, in a way that avoids the curse of dimensionality if the degree of anisotropic smoothness is sufficiently large (“$\mathcal{O}(d)$”), while leaving open the aspects of learning and optimization. The condition of belonging to a Barron’s space requires sufficient fast decaying rate for the Fourier transform of $\nabla F$, which is a condition that gets stronger with the dimension, except for very special cases, among which, notably the single- and multi-index models that we discuss momentarily, where the Fourier transform of $\nabla F$ is a singular measure. This requirement excludes intrinsically low-dimensional models from the Barron’s space since, as it is well-known in harmonic analysis, the Fourier transform of a measure supported on a general nonlinear curve decays slowly (Arkhipov et al., 1987; Brandolini et al., 2007). Similar arguments can be generalized to our main model, where the distribution is supported in a neighborhood of some nonlinear curve; our main model here is therefore typically not included in the Barron’s space, though both the Barron space and our main model include the single-index model as a special case.
@@
having some Hölder regularity, and $r\leq d$. For general function compositions, (Juditsky et al., 2009) proves that in some situations, when $g$ is sufficiently smooth, there is an improved min-max nonparametric rate of estimation of $F=f\circ g$, but still subject to the curse of dimensionality. Recent works combine anisotropic smoothness with composability, especially in the context of approximators or estimators constructed with deep neural networks. In this direction, (Schmidt-Hieber, 2020; Shen et al., 2021) consider spaces of functions that are compositions of low-dimensional functions with anisotropic smoothness conditions, studying the dependence of approximability and learning risk on the dimensionality of the spaces involved and the smoothness of the corresponding functions. Unfortunately the inner-most such function is defined on the original high-dimensional space $\mathbb{R}^{d}$, so in general these results do not circumvent the curse of dimensionality, unless again the (anisotropic, in general) regularity of such function scales with $d$. Another direction is to design different types of neural networks that exploit nonlinear compositions, see (Lai and Shen, 2021; Liu et al., 2024b). Most of these results only address whether a function can be well-approximated by a neural network, and do not cover the learning problem nor address computational aspects, in particular whether some optimization algorithm can find efficiently the parameters of the desired estimator. Another aspect that is often left unaddressed is the dependency of constants on the ambient dimension, which is unfortunately often exponential: see Shamir (2020) for a discussion of some of these aspects, and Shen et al. (2021) for bounds on such constants in some of these cases. In the parametric setting, models where $g$ is a polynomial and $f$ is a function in a known finite-dimensional space have been considered; for example Wang et al. (2024); Lee et al. (2024) (and related work referenced therein) consider special classes of polynomials $g$ for which a customized layer-by-layer training of a neural network leads to estimators of $F$ that are not cursed by the dimensionality.
@@
Single- and multi-index models. Classical examples of structural assumptions based on function composition for which the curse of dimensionality can provably be avoided include single- and multi-index models, as well as generalized linear models (Stone, 1982; Hastie and Tibshirani, 2014; Breiman and Friedman, 1985; Horowitz and Mammen, 2007). In the case when $g=G:\mathbb{R}^{d}\to\mathbb{R}^{r}$ is a linear operator, this is called the multi-index model, and it implies that the function $F\in\mathcal{C}^{s}(\mathbb{R}^{d})$ only depends on a small number $r$ of linear features: $F(x)=f(Gx)$ for some link function $f\in\mathcal{C}^{s}(\mathbb{R}^{r})$, matrix $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{r}$, and the projection of $X$ on the range of $G$ is sufficient for regression. The particular case $r=1$ is called the single-index model, where the function $F$ has the structure $F(x)=f(\langle v,x\rangle)$ for some unknown index vector $v\in\mathbb{S}^{d-1}$ and unknown link function $f\in\mathcal{C}^{s}(\mathbb{R}^{1})$. These models have been intensively studied: Stone (1982) conjectured that the min-max rate for regression for the single-index model is $n^{-\frac{s}{2s+1}}$ (resp. $n^{-\frac{s}{2s+r}}$ for multi-index models), coinciding with the one-dimensional min-max nonparametric rate, thereby escaping the curse of dimensionality. This rate is achieved with kernel estimators in (Hardle and Stoker, 1989, Theorem 3.3) and (Horowitz, 1998, Section 2.5), and the index $v$ can be estimated at the parametric rate $\mathcal{O}(n^{-\frac{1}{2}})$ under suitable assumptions. The existence of an estimator converging at rate $n^{-\frac{s}{2s+1}}$ is shown in (Györfi et al., 2002, Corollary 22.1), and (Gaïffas and Lecué, 2007, Theorem 2) demonstrates that $n^{-\frac{s}{2s+1}}$ is indeed the min-max rate. These models have also received attention recently both for their connections to “feature learning” in neural networks, see for example Lee et al. (2024) and references therein for learning single-index models with neural networks (trained in a suitable layer-wise fashion), as well as Radhakrishnan et al. (2024), and references therein, showing that deep neural networks appear to exploit low-dimensional linear subspaces for classification and prediction.
@@
Several methods for jointly estimating the index vector $v$ (or the matrix $G$) and $f$ were developed over the years, with varying tradeoffs between the assumptions needed to produce optimal or near-optimal estimators and the computational cost for constructing them, see Lanteri et al. (2022) for an extended discussion. One category includes semiparametric methods based on maximum likelihood estimation (Ichimura, 1993; Hardle et al., 1993; Delecroix et al., 2003; Delecroix and Hristache, 1999; Delecroix et al., 2006; R. J. Carroll and Wand, 1997) and M-estimators that produce $\sqrt{n}$-consistent index estimates under general assumptions, but with computationally demanding implementations, relying on sensitive bandwidth selections for kernel smoothing and on high-dimensional joint optimization of $f$ and $v$. Another category includes direct methods: for example Average Derivative Estimation (Stoker, 1986; Hardle and Stoker, 1989) estimates the index vector $v$ by exploiting its proportionality to $\nabla F$. Early implementations of this idea suffer from the curse of dimensionality due to the use of kernel estimation for the gradient. The technique of (Hristache et al., 2001) uses an iterative scheme to gradually adapt an elongated neighbourhood window. Gradient-based methods (stochastic and non-stochastic) have been studied in the context of neural networks (Lee et al., 2024).
@@
A category of techniques particularly relevant to the present work includes conditional, or inverse regression methods, which derive their estimators from statistics of the conditional distribution of the explanatory variable $X$ given observations of the (noisy) response variable $Y$. Prominent examples include sliced inverse regression (Duan and Li, 1991; Li, 1991), sliced average variance estimation (Cook, 2000), simple contour regression (Li et al., 2005; Coudret et al., 2014), with its analysis (in the multi-index case) in the work (Li and Wang, 2007), which yields an estimator with near-optimal rates for the multi-index model. Conditional methods partition the range of the response variable $Y$ into small intervals and consider their pre-images that are, when $f$ is monotone, slices with the thinnest side along the index direction $v$. They are often straightforward in implementation, consisting of the computation of conditional empirical moments, or other statistics related to the level sets of $f$, and can have only one parameter to tune: the width of slices. Unfortunately, several of the works above, including (Hristache et al., 2001; Li and Wang, 2007), while obtaining asymptotic learning rates that avoid the curse of dimensionality and are in some cases min-max optimal or near-optimal, require a minimal number of samples that is exponential in the ambient dimension $d$ (either explicitly, or through constants exponential in $d$ in the bounds).
@@
Besides the statistical view that the single- and multi-index model do not entail a statistical cost cursed by the ambient dimension, another critical problem is to find algorithms with reasonable computational cost to implement a statistically optimal estimator for these models. For example, even in the case of single-index models, $M$-estimators typically required high-dimensional non-convex optimization in $v$ and $f$; methods based on optimizing over $v$, even when combined with random sampling, may scale exponentially in $d$ due to the need of obtaining points inside a narrow cone around the unknown $v$. Conditional methods can often be implemented in a computationally efficient way, running in time $\mathcal{O}(C_{d}n)$ (up to $\log n$ factors, and with the exception of contour regression methods, which scale quadratically in $n$), with a constant $C_{d}$ a low-order polynomial in $d$. Lanteri et al. (2022) discuss in detail these techniques for the single-index model, and introduce of a variation, called Smallest Vector Regression, which enjoys optimal (possibly up to log terms) statistical guarantees up to log factors, provides a theoretically optimal choice for selecting of the slice width parameter (a crucial parameter whose choice is often not discussed (Coudret et al., 2014, p. 75), and is amenable to a computationally-efficient implementation with cost $\mathcal{O}(d^{2}n\log n)$, all without the curse of dimensionality in the exponents and in the constants, for the rate, the minimum sample requirements and the computational cost.
@@
In conclusion, while the situation is rather well-understood for compositional models $F=f\circ g$ with $g$ linear, much is open when $g$ is nonlinear, in which case it is not clear when the compositional structure allows one to circumvent the curse of dimensionality, except in cases where $g$ has high regularity (global or anisotropic), increasing with $d$, or (in particular) belongs to certain families of polynomials. In fact, we are not aware of models where $g$ is nonlinear, with regularity not infinite nor growing with $D$, for which $F$ may be estimated without the curse of dimensionality. We propose a simple model that, perhaps surprisingly, without strong regularity assumptions, still is amenable to nonparametric estimators that avoid the curse of dimensionality, both statistically and computationally.